The number of calls coming into an hotels reservation centre is poisson random variable with mean 3

i) Find the probability that no calls come in a given one minute period

ii)Assume that the number of calls arriving in two different minutes are independent. Find the probability that at least two calls will arrive in a given two minute period.


1
Expert's answer
2021-08-10T13:39:52-0400

i) Let "X=" the number of calls arriving  in a given one minute period: "X\sim Po(\lambda t)"

Given "\lambda=3, t=1"

"\lambda t=3(1)=3"


"P(X=0)=\dfrac{e^{-3}\cdot3^0}{0!}=e^{-3}\approx0.049787"


ii) Let "X=" the number of calls arriving in a given two minute period: "X\sim Po(\lambda t)"

Given "\lambda=3, t=2"

"\lambda t=3(2)=6"


"P(X\geq 2)=1-P(X=0)-P(X=1)"




"=1-\dfrac{e^{-6}\cdot6^0}{0!}-\dfrac{e^{-6}\cdot6^1}{1!}=1-7e^{-6}\approx0.982649"